### v1

# Agentic Design Review System

Abstract
Evaluating graphic designs involves assessing it from multiple facets like alignment, composition, aesthetics and color choices. Evaluating designs in a holistic way involves aggregating feedback from individual expert reviewers. Towards this, we propose an Agentic Design Review System (Agentic-DRS), where multiple agents collaboratively analyze a design, orchestrated by a meta-agent. A novel in-context exemplar selection approach based on graph matching and a unique prompt expansion method plays central role towards making each agent design aware. Towards evaluating this framework, we propose DRS-Bench benchmark. Thorough experimental evaluation against state-of-the-art baselines adapted to the problem setup, backed-up with critical ablation experiments brings out the efficacy of Agentic-DRS in evaluating graphic designs and generating actionable feedback. We hope that this work will attract attention to this pragmatic, yet under-explored research direction.
1 Introduction
Graphic designs like flyers, posters, invitation-cards, etc. are harmonious compositions of images, text, shapes and their colors, nicely laid-out aesthetically, to convey the meaning intended by their designer. They have become ubiquitous in our daily lives from the brochure of your new car to the birthday invitation of your toddler.
With the availability of do-it-yourself design tools, amateurs and novice designers are empowered to create professional designs. With the proliferation in the use of such designs in social media platforms, these tools are getting increasingly popular. Novice designers lack deep understanding of design principles like balance, emphasis, unity, white-space usage and so on, which would have a profound impact of their final generation. A tool that can automatically analyze a design and provide actionable feedback would be of immense value for such designers.
With the recent advancements in Multi-modal LLMs [42, 76] and Diffusion Models [56], researchers have introduced novel approaches [31, 28] for generating graphic designs from textual prompts. This is a significant advancement from the earlier efforts [27, 37, 46, 41] which generate just the layout (positioning information of elements), to generating the entire design, with the content filled in. As these technologies mature, they would truly enable Human-AI co-creation for design generation. A yard-stick to measure the progress of design generation methods would be a design-evaluator that introspects a design across multiple dimensions such as typographic quality, color consistency and semantic coherence.
Figure 1: Evaluating a graphic design involves assessing it across multiple dimensions like visual coherence, semantic grouping, typographic clarity and so on. Inspired by peer-review system in conferences, we propose to build the first Agentic framework for design evaluation and feedback generation. We propose a novel approach to infuse design knowledge into agents, and allow them to collaboratively review and access the input design in Sec. 3.
Judging whether a design is good or bad is hard and subjective as characteristics of designs are inherently tacit [^1] Dictionary meaning: understood or implied without being stated. [61]. Tacit principles like color pairing for a specific demography, is hard to objectively define and quantify making design evaluation extremely challenging. Heuristic approaches [51, 50, 49, 21, 5, 58, 73] try to mathematically quantify design characteristics like alignment, overlap and whitespace to give a score, but fail to see the overall global harmony of the designs. Learning based approaches [79, 57, 13, 65] treats the design evaluation as either a regression or classification task. Presence of a diverse enough dataset is a prerequisite for its performance. Recently, evaluations form GPT-4o [25] has found to correlate well with human judgment of graphic designs [20]. Building on this trend, we propose a holistic evaluation suite to access the quality of graphic designs, and generate actionable feedback for the designers.
By virtue of the huge amount of training data, and the learning paradigm, Multimodal LLMs (MLLMs) such as GPT-4o [25] posses novice-level awareness of the characteristics of a good graphic design. Our first contribution is to enhance the design awareness of these models using a novel graph matching based in-context exemplar selection approach (Sec. 3.1) and a structured description based prompt expansion strategy (Sec. 3.2). A single instance of such a design-aware MLLM would not be able to assess a design across the variety of dimensions on which graphic designs should be evaluated. Towards this end, we propose the first Agentic framework for design evaluation and feedback generation. Each agent will be specializing in a specific aspect of evaluation like color harmony, typographic quality, alignment consistency and so on. Given an input design, a meta-agent spawns-off these agents as necessary (could even dynamically control what aspect the agent should evaluate a design on, referred to as dynamic agents in Sec. 3), and collates the independent feedback after the review process to generate scores and actionable feedback.
Towards evaluating our approach, we introduce DRS-Bench, a holistic benchmark suite containing $15$ design attribute definitions, $4$ datasets, new evaluation metrics and strong baselines. Our experimental analysis showcases the effectiveness of using design-aware MLLMs in an Agentic framework for design evaluation and feedback generation.
Our key contributions are summarized below:
  • We introduce the first Agentic evaluation framework that can score designs and generate actionable feedback.
  • We enhance the design awareness of Multi-modal LLMs with a novel graph matching based exemplar selection approach and structured description based prompt expansion.
  • We introduce DRS-Bench, a holistic framework for assessing design evaluation quality and feedback generation.
  • Through rigorous experimentation, we bring out the efficacy of our proposed approach, clearly out-performing the state-of-the-art baselines adapted to the task.
2 Related Work
Design Evaluation. Design evaluation has been widely studied across domains and development stages. Early heuristic-based methods [4, 22, 50, 51, 72] struggled to capture semantic design attributes. Machine learning models [12] improved evaluation but lacked comparative scoring mechanisms, leading to siamese-based approaches [80, 16]. Kong et al. [34] trained a model for aesthetic assessment, while Tabata et al. [62] perturbed layouts to generate poor designs for scoring. Recent MLLM-based methods focus on design generation [30] or assess limited attributes [30, 20] but lack comprehensive, actionable feedback grounded in a comprehensive set of design attributes, an aspect we address.
MLLM Agents. Agentic workflows using large models are widely adopted for reasoning-based tasks. While early visual agents were fine-tuned for specific tasks [59], recent MLLMs enable broader agentic workflows [67]. These models integrate multi-modal encoders, allowing agents to tackle diverse tasks, including games [7, 69, 14], design [60, 36], human-computer interaction [81, 83, 33, 74], and audio understanding [77, 71, 24, 44]. These agents use structured planning and reasoning for complex problem-solving. However, to the best of our knowledge, no prior work has explored a multi-agent, multi-modal agentic workflow for Design Evaluation. We take the first step in this direction.
In-context Learning in MLLMs. In-context learning (ICL), where example inputs and outputs are included in prompts, is widely used in NLP to guide models toward relevant information for downstream tasks [11, 19, 23, 47, 48]. ICL has been extended to multi-modal applications [43, 75, 78], improving MLLM reasoning primarily for text-based tasks [2, 32, 35, 38, 39]. Jia et al. [29] demonstrated that visual context enhances ICL reasoning but within a fine-tuning framework. Training-free ICL sample selection has also been explored [55, 15], though limited to text [55] or global features [15]. In contrast, given our focus on design evaluation, we introduce a graph-based retrieval mechanism leveraging local features for in-context exemplar design retrieval while preserving fine-grained structural and semantic relationships among the various design elements.
3 Methodology
Figure 2: Overview of the proposed design evaluation pipeline. To systematically evaluate designs, GRAD constructs a graph representation of design elements using edge and node matching techniques, enabling structured retrieval of in-context design examples from a curated library (Sec 3.1). Structured Design Description (SDD) module generates design descriptions (Sec 3.2) to anchor the responses of each agent. The selected $K$ designs, query design along with its description inform the review process, where a meta agent coordinates static and dynamic agents to assess disparate design attributes - part of the Agentic Design Review System. Static agents focus on a fixed set of attributes (e.g., typography), whereas dynamic agents evaluate the attributes which are contextualized to specific designs (e.g., stylistic qualities). The meta agent consolidates these insights into a final rating and provides actionable feedback for design improvements (Sec 3.3).
Evaluating designs and generating actionable feedback are non-trivial and inherently complex tasks, requiring a nuanced understanding of aesthetics, functionality, user expectations, and intricacies of designs. The seminal work from Haraguchi et al. [20], shows how MLLMs such as GPT-4o are able to evaluate designs correlating with human ratings. This is an important first step towards evaluating designs in a systematic manner. Building on this, we further enhance the design capabilities of these MLLMs by making them more design-aware, in a training-free way, by introducing a novel graph-based in-context design selection approach (Sec 3.1) and anchoring MLLMs for designs with visually grounded structured descriptions (Sec 3.2). Given a query design to be evaluated, these key innovations assist in dynamically retrieving semantically and structurally relevant design examples from a design data library, thereby improving contextual understanding and enhancing the accuracy and robustness of design evaluation.
Scoring and generating feedback for a design should account for multiple aspects like visual appeal, semantic coherence, typographic quality, etc. Towards generating a holistic feedback that accounts for these multiple perspectives, we draw inspiration from how humans review their creative generations. The peer-review process in conferences solves a similar task by having a group of expert reviewers with diverse specializations (or expertise) analyzing the same subject from multiple perspectives, and meta-reviewers aggregating their opinions. This naturally inspires us to propose an Agentic Design Review System (as demonstrated in Fig 1), which to the best of our knowledge is a first effort to formalize such a collaborative review system for graphic design evaluation in Sec 3.3. By mirroring the principles of expert-driven decision-making, Agentic-DRS not only improves the robustness of design evaluations but also fosters a more explainable and structured critique process. This allows for nuanced feedback that balances subjective creativity with objective design principles, enabling designers to receive context-aware, multi-faceted insights that drive meaningful improvements in their work.
3.1 Graph-based Design Exemplar Selection
Conventional in-context example selection approaches often rely on global CLIP feature-based similarity, which may overlook finer design relationships such as spatial composition, alignment, grouping information, reading order, and contextual interplay among elements. Our GRAph-based Design exemplar selection method (GRAD) encodes semantic, spatial, and structural relationships by constructing localized representations of the graphic designs by offering certain advantages: (i) Preservation of structural and semantic relationships: unlike global feature matching , our graph representation captures relative positions and proximity of design elements. (ii) Adaptability to open-world designs: by computing similarities based on both semantic and spatial embeddings, the selection method generalizes across diverse design styles. To facilitate structured comparison and retrieval of reference designs, we define a query design (to be evaluated) $\mathcal{D}^{Q}$, and designs present in the In-context (IC) data library $\mathcal{D}^{\text{IC}}\in\mathcal{D}$ in the form of graphs $\mathcal{G}_{\mathcal{D}^{j}}=(\mathbb{V}^{j},\mathbb{E}^{j}),\;j\in\{Q,\text{IC}\}$, where $\mathbb{V}$ represents the vertices, and $\mathbb{E}$ denotes the edges of the graph, the superset of designs is $\mathcal{D}$. Graph creation is done in two ways:
(i) Presence of bounding boxes: in the presence of layout information metadata for designs, the bounding box (i.e., coordinate) information $\mathcal{BB}^{j}$ can be obtained for all the elements in a graphic design $\mathcal{D}^{j}$. The design components $V^{j}$ can be typically obtained by cropping the design with the coordinate information $\mathcal{BB}^{j}$: $V^{j}(i)=\text{Crop}(D^{j},\mathcal{BB}^{j}(i)),\;\forall i\in{1,...,N_{V^{j}}}$, $N_{V^{j}}$ being number of design elements in $\mathcal{D}^{j}$ ($j\in\{Q,\text{IC}\}$), $\text{Crop}(\cdot)$ being the cropping operation. Each such textual and visual elements are subsequently passed through CLIP to obtain respective set of embeddings:
$\phi(V^{j})(i)={\text{CLIP}(V^{j}(i))},\;\forall i\in{1,...,N_{V^{j}}}$ (1)
These embeddings of the corresponding design elements form the vertices $\mathbb{V}=\{\phi(V^{j})(i)\}_{i=1}^{N_{V}}$ of the graph. The edges are composed of spatial and semantic distances, where spatial distance $d_{\text{spatial}}(u,v)$ between two elements (i.e., nodes of a graph) $u$ and $v$ is computed based on the normalized L2 norm between the bounding box centroids $\mathcal{BB}^{j}(u)$ and $\mathcal{BB}^{j}(v)$ and the semantic distance $d_{\text{semantic}}(u,v)$ is computed as cosine distance between the respective embeddings $\mathbb{V}(u)$ and $\mathbb{V}(v)$. The combined edge weight is given as:
$d^{j}(u,v)=d_{\text{spatial}}^{j}(u,v)+d_{\text{semantic}}^{j}(u,v)$ (2)
(ii) Absence of bounding boxes: in the absence of layout information metadata for designs (for datasets without bounding-box details), we pass the design through a CLIP-ViT encoder and obtain the patch-level features from the last layers of corresponding visual transformer encoders. We subsequently use these encoded local-feature vectors to construct the graph where the vertices $\mathbb{V}$ contain the embedded features of patches and the edges $\mathbb{E}$ comprise of the semantic (cosine) distance between the vertices (patch-level embeddings). Since we do not have a clear structural information present (i.e., absence of boxes) we stick to using semantic relationships only.
Assuming a input design where inherently related elements are not grouped together, our approach GRAD will look for designs in the in-context library not only whose content is semantically close to the design under evaluation (via node matching) but also whose related elements are not clearly grouped (via edge matching). Therefore, GRAD enables structure-aware retrieval, improving the quality of selected in-context examples. This is achieved by the suitable combination of Wasserstein (node matching) [53, 9, 6, 45] and Gromov-Wasserstein (edge matching) distances as described below. Given a query design $\mathcal{D}^{Q}$, we select the most relevant in-context examples $\{\mathcal{D}_{k}^{\text{IC}}\}_{k=1}^{K}$ from a dataset $\mathcal{D}^{\text{IC}}$ using Wasserstein and Gromov-Wasserstein distances.
Wasserstein Distance (WD) calculates the pairwise distances between the node embedding sets $\mathbb{X}^{Q}$ and $\mathbb{Y}^{k}$ of the respective graphs $\mathcal{G}_{\mathcal{D}^{Q}}$ and $\mathcal{G}_{\mathcal{D}^{\text{IC}}_{k}}$. Considering two discrete distributions, $\psi^{Q}\in\mathbf{P}(\mathbb{X}^{Q})$ and $\psi^{k}\in\mathbf{P}(\mathbb{Y}^{k})$, where $\psi^{Q}=\sum_{i=1}^{n}z^{Q}(i)\delta_{x^{Q}}(i)$ and $\psi^{k}=\sum_{j=1}^{m}z^{k}(j)\delta_{y^{k}}(j)$; $\sum_{i}z^{Q}(i)=1=\sum_{j}z^{k}(j)$ where $z^{Q}$ and $z^{k}$ are the respective weight vectors for the probability distributions $\psi^{Q}$ and $\psi^{k}$; and $\delta_{x^{Q}}$ is the Dirac-delta function centered on support point $x^{Q}$ in the embedding space. The WD distance between $\psi^{Q}$ and $\psi^{k}$ is defined as:
$\mathbb{D}_{\text{W}}(\mathcal{G}_{\mathcal{D}^{Q}},\mathcal{G}_{\mathcal{D}^{\text{IC}}_{k}})\triangleq\min_{\mathbf{\Phi}\in\mathbf{\Theta}(z^{Q},z^{k})}\sum_{i,j}\mathbf{\Phi}_{ij}\cdot c(x_{i},y_{j})$ (3)
where $\mathbf{\Theta}(z^{Q},z^{k})=\{\mathbf{\Phi}\in\mathbb{R}_{+}^{n\times m}|\mathbf{\Phi}\mathbf{1}_{m}=z^{Q},\mathbf{\Phi}^{\top}\mathbf{1}_{n}=z^{k}\}$, $c(x^{Q}(i),y^{k}(j))$ is cosine distance similarity metric, and $\mathbf{\Phi}$ is the transport plan, interpreting the amount of mass shifted from distributions $\psi^{Q}(i)$ to $\psi^{k}(j)$. An exact solution to the above expression leads to a sparse representation of transport plan $\mathbf{\Phi}$ with at most $(2\cdot\text{max}(m,n)-1)$ non-zero elements, ensuring an explainable and robust retrieval [10]. A detailed set of steps is provided in Supp Algorithm 1.
Gromov-Wasserstein Distance (GWD) helps in matching the edges of the graphs and preserves graph topology by computing distances between pairs of nodes thereby ensuring inter-graph structural alignment [52, 3]. In the same discrete graph matching setting, GWD can be mathematically represented:
$\scriptsize\mathbb{D}_{\mathrm{GW}}(\mathcal{G}_{\mathcal{D}^{Q}},\mathcal{G}_{\mathcal{D}^{\text{IC}}_{k}})\triangleq\min_{\hat{\mathbf{\Phi}}\in\mathbf{\Theta}(z^{Q},z^{k})}\sum_{i,i^{\prime},j,j^{\prime}}\hat{\mathbf{\Phi}}_{ij}\hat{\mathbf{\Phi}}_{i^{\prime}j^{\prime}}\mathcal{L}\left(x^{Q}_{i},y^{k}_{j},x^{Q}_{i^{\prime}},y^{k}_{j^{\prime}}\right)$ (4)
where inter-graph structural similarity between two node pairs $(x^{Q}_{i},x^{Q}_{i^{\prime}})$ and $(y^{k}_{j},y^{k}_{j^{\prime}})$ is represented as $\mathcal{L}\left(x^{Q}_{i},y^{k}_{j},x^{Q}_{i^{\prime}},y^{k}_{j^{\prime}}\right)=\|d^{Q}(x^{Q}_{i},x^{Q}_{i^{\prime}})-d^{k}(y^{k}_{j},y^{k}_{j^{\prime}})\|$, $d^{Q}$ being the edge weight between a node pair in graph $\mathcal{G}_{\mathcal{D}^{Q}}$ (see Equation 2). Transport plan $\hat{\mathbf{\Phi}}$ is periodically updated to align the edges in different graphs belonging to disparate designs. A detailed set of steps is provided in Supp Algo 2.
Algorithm 1 Overall Pipeline
1:Query Design: $\mathcal{D}^{Q}$, In-context Designs: $\mathcal{D}^{\text{IC}}$, In-context samples to be retrieved: $K$, Static Attribute Buckets: $B_{A_{S}}$, Meta Agent: $A_{M}$, Static Agents: $A_{S}$, Number of Static Agents: $N_{S}$, World Attributes: $\mathcal{W}$, Bounding Boxes: $\mathcal{BB}$, CLIP Encoder: $\text{CLIP}(\cdot)$ Graph Construction Module: $\Omega(\cdot)$, Design Description module: $\Psi(\cdot)$, Wasserstein Distance func.: $\mathbb{D}_{\text{W}}(\cdot)$, Gromov-Wasserstein Distance function: $\mathbb{D}_{\text{GW}}(\cdot)$, Balancing factors: $\alpha$, $\lambda$, Index Sorting function: Argsort($\cdot$), List Element Insertion function: Insert($\cdot$).
2:Dynamic Attribute Buckets: $B_{A_{D}}$, Dynamic Agents: $A_{D}$, Number of Dynamic Agents: $N_{D}$, Feedback: $\mathcal{F}$, Attribute Rating: $\mathcal{R}$, Query Graph: $\mathcal{G}_{\mathcal{D}^{Q}}$, $k^{\text{th}}$ in-context Graph: $\mathcal{G}_{\mathcal{D}^{\text{IC}}_{k}}$, Graph Dissimilarity Scores List: $\mathbb{S}_{L}$, Design Description: $\mathbb{T}$.
[…]
4:for $\mathcal{D}^{\text{IC}}_{k}\in\mathcal{D}^{\text{IC}}$ do
5:  $\mathcal{G}_{\mathcal{D}^{\text{IC}}_{k}}\leftarrow\Omega(\mathcal{D}^{\text{IC}}_{k})$ $\triangleright$ graph construction, Sec 3.1
6:  $\mathbb{S}_{\text{l}}\leftarrow\alpha\mathbb{D}_{\text{W}}(\mathcal{G}_{\mathcal{D}^{Q}},\mathcal{G}_{\mathcal{D}^{\text{IC}}_{k}})+(1-\alpha)\mathbb{D}_{\mathrm{GW}}(\mathcal{G}_{\mathcal{D}^{Q}},\mathcal{G}_{\mathcal{D}^{\text{IC}}_{k}})$ $\triangleright$ Sec 3.1
7:  $\mathbb{S}_{\text{g}}\leftarrow 1-\texttt{cos}(\text{CLIP}(\mathcal{D}^{Q}),\text{CLIP}(\mathcal{D}^{\text{IC}}_{k}))$ $\triangleright$ Sec 3.1
8:  $\mathbb{S}(\mathcal{D}^{Q},\mathcal{D}^{\text{IC}}_{k})\leftarrow\mathbb{S}_{\text{l}}+\mathbb{S}_{\text{g}}$ $\triangleright$ Sec 3.1
9:  $\mathbb{S}_{L}\leftarrow\texttt{{Insert}}(\mathbb{S}_{L},\mathbb{S}(\mathcal{D}^{Q},\mathcal{D}^{\text{IC}}_{k}))$ $\triangleright$ inserting scores to list
[…]
14:for $n_{s}\in{1,\cdots,N_{S}}$ do $\triangleright$ reviewing, Sec 3.3
15:  $\mathcal{R}_{S}(n_{s}),\mathcal{F}_{S}(n_{s})\leftarrow A_{S}^{(n_{s})}(\mathcal{D}^{Q},\mathbb{T}_{\mathcal{D}^{Q}},\mathcal{D}^{\text{IC}}[\mathbb{I}_{K}],B_{A_{S}}^{(n_{s})})$
16:for $n_{s}\in{1,...,N_{D}}$ do $\triangleright$ reviewing, Sec 3.3
17:  $\mathcal{R}_{D}(n_{d}),\mathcal{F}_{D}(n_{d})\leftarrow A_{D}^{(n_{d})}(\mathcal{D}^{Q},\mathbb{T}_{\mathcal{D}^{Q}},\mathcal{D}^{\text{IC}}[\mathbb{I}_{K}],B_{A_{D}}^{(n_{d})})$
[…]
19:return $\mathcal{R},\mathcal{F}$
The combined dissimilarity (since we are computing distances) score $\mathbb{S}_{\text{l}}$ is a weighted combination of Eqs. 3 and 4, based on which top-$K$ designs are selected from the in-context design library $\mathcal{D}^{\text{IC}}\subset\mathcal{D}$.
$\mathbb{S}_{\text{l}}=\alpha\mathbb{D}_{\text{W}}(\mathcal{G}_{\mathcal{D}^{Q}},\mathcal{G}_{\mathcal{D}^{\text{IC}}_{k}})+(1-\alpha)\mathbb{D}_{\mathrm{GW}}(\mathcal{G}_{\mathcal{D}^{Q}},\mathcal{G}_{\mathcal{D}^{\text{IC}}_{k}})$ (5)
Along with $\mathbb{S}_{\text{l}}$ (on local representations), we add global scores on renditions: $\mathbb{S}_{\text{g}}=1-\texttt{cos}(\text{CLIP}(\mathcal{D}^{Q}),\text{CLIP}(\mathcal{D}^{\text{IC}}_{k}))$ where cos($\cdot$) is the cosine similarity function. The final expression becomes: $\mathbb{S}(\mathcal{D}^{Q},\mathcal{D}^{\text{IC}}_{k})=\mathbb{S}_{\text{l}}+\mathbb{S}_{\text{g}}$.
3.2 Structured Design Description (SDD)
For the input query design $\mathcal{D}^{Q}$, our goal is to generate textual descriptions $\mathbb{T}_{\mathcal{D}^{Q}}$ which contains description of elements (images, icons, texts, etc.) and how they are structured hierarchically. For example, a textual description may look like: “A title ‘ABC’ at the top $bb_{11}^{Q},bb_{12}^{Q},bb_{13}^{Q},bb_{14}^{Q}$] with an image of ‘X’ below it $bb_{21}^{Q},bb_{22}^{Q},bb_{23}^{Q},bb_{24}^{Q}$]. Below the image, there is a text contaning ‘DEF’, … ”. Passing both the graphic design and a textual description with bounding box ($\mathcal{BB}^{Q}=\{bb_{ij}^{Q}\}$) information (optional, as available from the layout metadata) improves design attribute understanding and anomaly detection by combining visual perception with explicit structural and semantic details along with element relationships. Such textual descriptions are superior to merely feeding the raw metadata information (e.g., xml or json) and assists in anchoring the MLLM responses in a detailed visual description. In addition, it makes design attribute understanding more robust to diverse layouts while facilitating clearer, more actionable feedback, reducing hallucinations. Notably, we obtain the descriptions by passing the designs through an MLLM $\Psi(\cdot)$ with suitable prompt.
$\mathbb{T}_{\mathcal{D}^{Q}}=\Psi\left(\mathcal{D}^{Q},\mathcal{BB}^{Q}\right);\text{if}\;\mathcal{BB}^{Q}\neq\emptyset,\;\text{else}\;\Psi\left(\mathcal{D}^{Q}\right)$ (6)
Method Discrete Evaluation (Classification) Continuous Evaluation (Correlation)
Afixa Infographic IDD GDE [20]
Acc $\uparrow$ Sens $\uparrow$ Spec $\uparrow$ Acc $\uparrow$ Sens $\uparrow$ Spec $\uparrow$ Acc $\uparrow$ Sens $\uparrow$ Spec $\uparrow$ Alignment $\uparrow$ Overlap $\uparrow$ Whitespace $\uparrow$
Heuristic-based Evaluation - - - - - - - - - 0.310 0.476 0.233
Gemini-1.5-Pro 59.45 62.11 60.18 54.88 55.61 54.85 64.37 67.89 65.59 0.586 0.759 0.641
Gemini-1.5-Pro + GRAD 62.19 64.08 62.45 56.21 60.72 56.18 68.21 68.65 67.16 0.623 0.778 0.676
Gemini-1.5-Pro + GRAD + SDD 65.62 67.95 66.31 59.76 63.84 59.05 69.09 68.94 70.42 0.671 0.783 0.691
Agentic-DRS $_{\text{Gemini-1.5-Pro}}$ 72.17 74.94 70.85 65.97 67.41 68.58 75.43 75.31 76.22 0.712 0.821 0.739
$\Delta_{\text{{Agentic-DRS}- Gemini-1.5-Pro}}$ 12.72 $\uparrow$ 12.83 $\uparrow$ 10.67 $\uparrow$ 11.09 $\uparrow$ 11.80 $\uparrow$ 13.73 $\uparrow$ 11.06 $\uparrow$ 7.42 $\uparrow$ 10.63 $\uparrow$ 0.126 $\uparrow$ 0.062 $\uparrow$ 0.098 $\uparrow$
GPT-4o 62.91 65.42 64.26 58.26 61.92 56.74 65.72 65.38 66.57 0.597 0.782 0.665
GPT-4o + GRAD 64.57 68.65 65.18 60.41 63.57 59.66 68.51 67.26 70.85 0.639 0.796 0.688
GPT-4o + GRAD + SDD 67.33 69.60 68.21 64.95 66.21 62.12 70.16 69.44 73.92 0.677 0.809 0.703
Agentic-DRS $_{\text{GPT-4o}}$ 75.29 77.65 72.53 69.53 75.37 71.94 76.78 74.56 80.31 0.722 0.834 0.748
$\Delta_{\text{{Agentic-DRS}- GPT-4o}}$ 12.38 $\uparrow$ 12.23 $\uparrow$ 8.27 $\uparrow$ 11.27 $\uparrow$ 13.45 $\uparrow$ 15.20 $\uparrow$ 12.28 $\uparrow$ 9.18 $\uparrow$ 13.74 $\uparrow$ 0.125 $\uparrow$ 0.052 $\uparrow$ 0.083 $\uparrow$
Table 1: Performance of Agentic-DRS on the DRS-Bench. Agentic-DRS outperforms baseline methods by substantial margins across all metrics in both evaluation protocols (discrete for attribute classification, and continuous for correlation with human labels).
3.3 Agentic-Design Review System
When evaluating a design, insights from design literature [17, 8, 66] highlight the critical importance of attributes such as alignment, overlap and spacing. Any effective review system must rigorously assess these foundational principles to ensure high-quality design evaluation. To achieve this, we introduce a structured, agentic review framework Agentic-DRS where specialized agents focus on specific design attributes. Consequently, we introduce a set of Static Agents (evaluators/reviewers), which are predefined and operate with fixed (static) roles, each dedicated to assessing well-established design attributes that are universally relevant across designs — such as alignment, overlap and spacing. These agents serve as the backbone of the evaluation system, ensuring that essential design attributes are consistently reviewed.
However, design evaluation is not solely a rule-based exercise; context-dependent attributes often influence how a design is perceived and interpreted. To capture these contextual nuances, we introduce a novel set of Dynamic Agents (evaluators/reviewers), which adapt based on the unique characteristics of a given design. These agents assess factors such as relative spacing, grouping, semantic effectiveness of communication, stylistic coherence, etc. which vary across different designs (i.e., dynamic in nature and contextualized on designs) and are not universally predefined.
The informed decision of which static and dynamic agents should be activated for a particular query design is exclusively handled by a Meta Agent (evaluators/reviewers), which intelligently plans the evaluation process, ensuring that the most relevant aspects of a design are meticulously scrutinized - this constitutes the planning phase. Once individual agents have assessed the design and provided their ratings (reviewing phase), the system must aggregate their insights into a coherent and actionable review. A consolidation mechanism, also managed by the Meta Agent, performs summarization, synthesizing feedback, resolving potential inconsistencies, and generating a unified evaluation report. This ensures that the final assessment is not just a collection of disjointed scores but a holistic, constructive critique that can guide meaningful improvements in the design. Notably, as opposed to prior design evaluation systems, either relying on predefined heuristics [51] or end-to-end learned scoring functions [16], both of which lack adaptability, our framework introduces an adaptive mechanism, (akin to a peer-review process), making our approach suitable for evolving design principles and diverse stylistic elements.
We formally define the agents participating in the design evaluation process, comprising of Meta Agent $A_{M}$, Static Agents $A_{S_{i}}\forall i\in\{1,2,...,N_{S}\}$, and Dynamic Agents $A_{D_{i}}\forall i\in\{1,2,...,N_{D}\}$. $\mathcal{S}$ represents the state space describing the design under evaluation. The joint action space is denoted by $\mathbb{A}$, where each agent $A_{i}$ executes actions $a_{i}$ based on its policy $\pi_{i}$. The transition function $\mathcal{T}:\mathcal{S}\times\mathbb{A}\to\mathcal{S}$ models the evolution of the design evaluation state as agents contribute their assessments. For a given query design $\mathcal{D}^{Q}$, the agents collaborate in a structured sequence, via a directed interaction network (Fig 2) which can be decomposed into the aforementioned 3 phases (planning, reviewing, summarization), which we describe below.
(i) Planning: The meta-agent $A_{M}$ plans the evaluation process by acting as a router which initiates the process by assigning static reviewers $A_{S_{i}}$ based on predefined criteria and dynamic reviewers $A_{D_{i}}$ based on attributes sampled from the open-world design principles (attributes) $\mathcal{W}$ which are deemed contextual and relevant for the design to be evaluated. Notably, each static evaluator $A_{S_{i}}$ is responsible for evaluating a specific set of design attributes with predefined buckets as $B_{A_{S}}(i)\triangleq\{w_{A_{S}}^{(i)}(1),...,w_{A_{S}}^{(i)}(k)\},\;w_{A_{S}}^{(i)}(k)\in W_{A_{S}}\subset\mathcal{W},\;i\in\{1,...,N_{S}\}$, which are embedded in their respective prompts $p_{A_{S}}^{(i)}\left(B_{A_{S}}(i)\right)\in\mathcal{P}$ (the collective prompt space is represented as $\mathcal{P}$). Whereas, for each dynamic evaluator, attributes are first sampled $W_{A_{D}}\subseteq\mathcal{W}-W_{A_{S}}$ and dynamically bucketed on the fly (as decided by the meta agent) into $N_{D}$ buckets where each $B_{A_{D}}(i)\triangleq\{w_{A_{D}}^{(i)}(1),...,w_{A_{S}}^{(i)}(l)\},\;w_{A_{D}}^{(i)}(l)\in W_{A_{D}},\;i\in\{1,...,N_{D}\}$ followed by dynamic prompt creation $p_{A_{D}}^{(i)}\in\mathcal{P}$.
$a_{M}\sim\pi_{M}(\cdot|s,\mathcal{D}^{Q},\mathcal{D}^{\text{IC}}_{K}),\quad\;s\sim\mathcal{S}$ (7)
(ii) Reviewing: Both groups of agents (static and dynamic) follow a policy $\pi_{i}$, determining its output based on the previous agent (i.e., meta agent), and the top-$K$ n-context (IC) samples $\mathcal{D}^{\text{IC}}_{K}$ from the in-context Data Library $\mathcal{D}^{\text{IC}}$ ($\mathcal{D}^{\text{IC}}_{K}\subseteq\mathcal{D}^{\text{IC}}\subset\mathcal{D}$), in the interaction network:
$a_{i}\sim\pi_{i}(\cdot|s,\{a_{M}\}_{A_{M}},\mathcal{D}^{Q},\mathcal{D}^{\text{IC}}_{K}),\quad s\sim\mathcal{S}$ (8)
where $A_{M}$’s outputs influence $A_{i}$. The aggregated actions $a=(a_{1},\dots,a_{N})$ collectively determine the next state: $s_{t+1}=\mathcal{T}(s_{t},a_{t})=\text{Concat}(s_{t},a_{t})$, where Concat represents the concatenation operation that updates the design evaluation context. These agents assign quantitative measures (design attribute ratings) and qualitative measures (actionable feedback), respectively denoted by $\mathcal{R}_{S}$ and $\mathcal{F}_{S}$ (for static agents), and $\mathcal{R}_{D}$ and $\mathcal{F}_{D}$ (for dynamic agents).
(iii) Summarization: The meta-agent $A_{M}$ collates the scores and integrates the feedback it receives from all static and dynamic reviewers to construct a final evaluation.
$a_{M}\sim\pi_{M}(\cdot|s,\{a_{j}\}_{A_{j}\in\text{Pred}(A_{M})}),\quad s\sim\mathcal{S}$ (9)
where $\text{Pred}(A(i))$ represents predecessor agent(s) who directly influence the outputs of the successor agent(s). Using this mechanism, redundant feedback is removed, refining the assessment and obtaining a final list of actionable feedback $\mathcal{F}$ (from $\mathcal{F}_{S}$, $\mathcal{F}_{D}$) and Attribute Ratings $\mathcal{R}$ (from $\mathcal{R}_{S}$, $\mathcal{R}_{D}$).
3.4 Overall Framework
We summarize the overall flow of our pipeline in Algo 1. Our key novelties are: (i) to adapt the graph matching algorithm based on localized representations (Supp Algos 1 - 2) for selecting the top-$K$ designs based on scores (lines 1 - 9, Algorithm 1), (ii) to anchor MLLM responses via structured design descriptions generated using design renditions and bounding boxes (line 10, Algo 1), and (iii) to introduce agentic design review framework comprising of meta, static and dynamic agents and involving planning, reviewing and summarization mechanisms (lines 11 - 16, Algo 1).
Method Metric GDE Infographic Afixa IDD
Agentic-DRS $_{\text{Gemini}}$ Human 0.744 0.682 0.720 0.742
[…]
Agentic-DRS $_{\text{GPT-4o}}$ Human 0.762 0.708 0.736 0.740
[…]
Table 2: Feedback evaluation of Agentic-DRS on both GPT-4o and Gemini versions. Strong correspondences can be observed across different variations of AIM and also with the human ratings.
4 Experiments and Results
4.1 DRS-Bench: Design Evaluation Benchmark
Graphic design principles provide essential guidelines for creating clear, coherent, and usable compositions [8, 17, 66]. Existing heuristic-based evaluation methods [51] often assess only a few principles, such as alignment, overlap, and white space, or provide a single goodness-of-fit score [16], overlooking key aspects like visual hierarchy, color pairings, and font compatibility. Despite advancements in automated design generation [46, 26, 68, 70, 63, 41, 18], there is no standardized evaluation framework or benchmark to capture critical design attributes. This gap hinders objective assessment, limits comparisons, and makes scaling reliable, data-driven design models challenging. Thus, we propose DRS-Bench, a unified benchmark for evaluating design effectiveness, enabling fair comparisons, improving automated design tools, and enhancing overall design quality.
Attributes. A core aspect of our benchmark is the multi-dimensional evaluation of designs through a comprehensive set of design attributes, collectively termed “World Attributes” ($\mathcal{W}$). These attributes capture fundamental design flaws and best practices. Text-rendering quality assesses legibility based on font size, weight, and clarity. Too many words penalizes cluttered text, while too many fonts and bad typography colors flag inconsistent styling and poor contrast. Composition and layout, alignment, and spacing ensure visual balance, favoring structured designs with clear margins and uniform spacing. Color harmony and wrong color palettes measure the effectiveness of color choices in maintaining aesthetic appeal. Style evaluates stylistic aspects and consistency, while grouping and image-text alignment address content organization and readability. Aesthetics considers overall visual appeal, penalizing overlap, bad images, and poorly arranged elements.
Datasets. In DRS-Bench, we include 4 datasets which are:
GDE: DRS-Bench leverages publicly available GDE dataset [20], hosting a large collection of 700 banner and poster designs each of which containing 3 attributes: alignment, overlap, and white space, on a 1 to 10 scale. Layout metadata information are not publicly available for GDE.
Afixa: It consists of 71 designs, in DRS-Bench as collected from a public platform Roboflow[1] which consists of yes or no values for 5 design attributes: wrong color palettes pairings, bad typo colors, bad images, too many words, too many fonts. Layout metadata information are not available.
Infographic: It consists of 55 samples collected from the Roboflow platform. This consists of the layout metadata information (xml) for the elements (i.e., bounding boxes for the elements). Each design has yes or no responses for all the 15 attributes (see Sec 4.1).
Internal Design Dataset (IDD): We internally collect 137 design samples each of which has yes or no responses to the 15 design attributes (see Sec 4.1). The designs are professionally curated and typically comprises of flyers, invitations, posters, albeit with some inconsistencies in design attributes. IDD consists of the layout metadata information (xml) with bounding boxes for the design elements.
4.2 Evaluation Metrics
Attribute Evaluation: Design flaws often coexist, impacting readability, style, and aesthetics. Assigning multiple labels allows for a more detailed evaluation rather than reducing it to a single goodness-of-fit score. For Afixa, Infographic, and IDD, attributes are labeled as yes or no, making it a multi-label classification problem. We report the mean Sensitivity, Specificity, and Accuracy, termed Discrete Evaluation. For GDE, attributes are rated on a continuous scale (1–10), so we follow [20] and report Pearson correlation with human labels, termed Continuous Evaluation.
[…]
4.3 Baselines
To the best of our knowledge, there does not exist any design-aware MLLMs let alone agentic frameworks. Hence, to compare Agentic-DRS, we introduce some baselines with two powerful MLLMs, Gemini-1.5 Pro [64] and GPT-4o [25], and subsequently adding design-aware components such as GRAD and SDD. However, these baselines are exclusively restricted to a single MLLM agent. Furthermore, notably, vanilla GPT-4o implementation for design has been proposed in [20] which we adopt in Table 1. Along with these, we also report heuristic evaluation results following [51] on the GDE dataset [20] only. Afixa, Infographic and IDD have design attributes which go well beyond conventional heuristics (e.g., alignment, overlap and spacing) and therefore cannot be evaluated with the heuristics-based method.
Figure 3: Generated feedback along with the design attributes which are found to be inconsistent for the input design evaluated (best viewed when zoomed). Feedback evaluation scores are also reported (as outlined in Sec 4.2).
In-context Sampling Method GDE (Avg.) Infographic (Acc) Afixa (Acc)
Random Selection (9) 0.743 63.75 72.06
Global Features (6) 0.752 64.39 72.94
Description-based (6) 0.755 67.05 73.98
GRAD (w/o Global Features) (5) 0.760 67.13 74.81
GRAD (4) 0.768 69.53 75.29
Table 3: Ablation on different in-context design selection methods for Agentic-DRS $_{\text{GPT-4o}}$. Best results are obtained with GRAD, preserving semantic, spatial and structural information. Best $K$ for top-$K$ retrieval are provided in brackets beside method.
GRAD SDD Infographic (Acc) IDD (Acc)
w/ Bnd. Box w/o Bnd. Box
63.11 69.75
64.52 71.03
65.71 72.86
66.48 73.95
67.82 74.57
69.53 76.78
Table 4: Effect of different components of Agentic-DRS $_{\text{GPT-4o}}$. Results on dataset with layout metadata information are reported. Substantial improvements are found upon adding GRAD and SDD.
$\alpha$ Infog. (Acc) Afixa (Acc)
0 68.10 73.95
1 66.85 72.86
0.25 68.94 74.38
0.75 67.21 73.14
0.5 69.53 75.29
Table 5: Ablation on $\alpha$ for Agentic-DRS $_{\text{GPT-4o}}$ on Infographic and Afixa datasets.
[…]
Figure 4: Impact of $K$ in GRAD for Agentic-DRS.
4.4 Results
[…]
Fig. 3 presents a qualitative example from the Infographics dataset, showcasing how Agentic-DRS generates actionable feedback for design evaluation. The feedback quality is validated by both AIM scores (Sec 4.2) and human ratings, normalized between 0 and 1 for fair comparison. Our GRAD module selects relevant samples for design understanding, while SDD grounds element relative positions, enhancing evaluation insights. This enables references to specific design elements (e.g., inconsistent spacing in the ”Speed Limit” section), improving feedback actionability. Additionally, stylistic aspects like elegance and minimality are correctly assessed by dynamic experts, spawned appropriately by the meta agent.
5 Discussions and Analysis
5.1 Analyzing the Impact of GRAD
Table 3 compares GRAD with four retrieval approaches. A random selection of exemplar designs yields the lowest scores on DRS-Bench. Using CLIP global features improves performance over random selection, consistent with prior findings [15]. A description-based approach, which retrieves top-$K$ samples using SDD-generated textual descriptions (Sec 3.2), outperforms global features, especially when design element descriptions with coordinates are available (e.g., Infographic). Removing global features from GRAD (Sec 3.1) leads to a slight performance drop, highlighting the need for both structural and global features.
GRAD achieves the best performance while requiring fewer in-context samples, as it preserves spatial and compositional relationships rather than treating designs as unordered features. Figure 4 shows similar trends in Accuracy across different $K$ values for GPT-4o and Gemini on the Infographic dataset, demonstrating GRAD’s robustness. Finally, Table 4 shows that even without SDD, GRAD significantly boosts performance, as seen in Table 1 for single MLLMs.
5.2 Assessing the Role of SDD
Design renditions capture aesthetic and spatial patterns, while textual descriptions provide an anchor for understanding. Table 4 shows that even without GRAD, SDD significantly boosts evaluation performance in Agentic-DRS, with similar trends in Table 1 for single MLLM models (GPT/Gemini). Performance improves further with bounding boxes, demonstrating that detailed visual descriptions with coordinate references (akin to phrase grounding or referring expressions) enhance model comprehension and anomaly detection in designs. We use structured descriptions over raw XML metadata because LLMs inherently process textual information more effectively.
5.3 Analyzing the Effect of $\alpha$
Table 5 shows the impact of different $\alpha$ values in balancing Wasserstein (WD) and Gromov-Wasserstein distances (GWD). Using only GWD ($\alpha=0$) outperforms WD alone ($\alpha=1$), but the best performance occurs at $\alpha=0.5$, denoting the importance of combining semantic feature matching with structural similarities for improved in-context retrieval.
6 Conclusion
We propose the first Agentic framework for critiquing a graphic design across multiple dimensions to generate scores and actionable feedback. This can serve as a tool in empowering novice designers to enhance their designs, and for evaluating the quality of designs generated by generative models [31, 28] alike. Our novel exemplar selection approach and prompt expansion technique is critical towards developing the framework, as validated through our extensive experimental analysis. A key next step would be to extend the framework to automatically applying the generated actionable feedback to the input design, thereby enabling a self-improving framework for graphic design generation.
Appendix
Appendix A Algorithms for Wasserstein and GW Distances
We provide the algorithms for Wasserstein and Gromov Wasserstein Distance computation respectively in Algorithhms 2 and 3 following [9, 52, 3, 53].
Algorithm 2 Wasserstein Distance Computation in GRAD
1:Initial Transport Plan: $\mathbf{\Phi}^{(1)}=\mathbf{1}\mathbf{1}^{\top}$; Initial scaled unity matrix: $\boldsymbol{\sigma}=\frac{1}{m}\mathbf{1_{m}}$; Cost (similarity) matrix b/w node pairs of two graphs: $\mathbf{C}_{ij}$; Total Optimal Transport steps: $\mathcal{N}_{\mathbf{\Phi}}$; Cost matrix decay factor: $\beta$; Scaled cost matrix: $\mathbf{\Upsilon}_{ij}={\rm e}^{-\frac{\mathbf{C}_{ij}}{\beta}}$.
2:Optimal Transport Plan: $\mathbf{\Phi}$; Wasserstein Distance: $\mathbb{D}_{\text{W}}$.
3:for $t\in\{1,2,3,\cdots\mathcal{N}_{\mathbf{\Phi}}\}$ do
4:  $\mathbf{Q}\leftarrow\mathbf{\Upsilon}\odot\mathbf{\Phi}^{(t)}$ $\triangleright$ $\odot$ is Hadamard product
5:  for $l\in\{1,\cdots,L\}$ do
6:   $\boldsymbol{\delta}\leftarrow\frac{1}{m\mathbf{Q}{\boldsymbol{\sigma}}}$, $\boldsymbol{\sigma}\leftarrow\frac{1}{m\mathbf{Q}^{\top}\boldsymbol{\delta}}$
7:  $\mathbf{\Phi}^{(t+1)}\leftarrow\text{diag}(\boldsymbol{\delta})\mathbf{Q}\text{diag}(\boldsymbol{\sigma})$
[…]
9:return $\mathbf{\Phi}$, $\mathbb{D}_{\text{W}}$
Algorithm 3 GW Distance Computation in GRAD
1:Nodes of Graphs $\mathcal{G}_{x},\mathcal{G}_{y}$: $\{x_{i}\}_{i=1}^{m},\{y_{j}\}_{j=1}^{n}$; Intra-domain Cost Matrices: $\mathbf{C}_{x}]_{i,i^{\prime}}=d(x_{i},x_{i^{\prime}}),[\mathbf{C}_{y}]_{j,j^{\prime}}=d(y_{j},y_{j^{\prime}})$; probability vectors: $p=\frac{1}{m}\mathbf{1_{m}},q=\frac{1}{n}\mathbf{1_{n}}$; Total Optimal Transport steps: $\mathcal{N}_{\hat{\mathbf{\Phi}}}$.
2:Optimal Transport Plan: $\hat{\mathbf{\Phi}}$; GW Distance: $\mathbb{D}_{\text{GW}}$.
[…]
4:for $t\in\{1,\cdots$$\mathcal{N}_{\hat{\mathbf{\Phi}}}$do
5:  $\mathcal{L}=\mathbf{C}_{xy}-2\mathbf{C}_{x}\hat{\mathbf{\Phi}}^{(t)}\mathbf{C}_{y}^{\top}$
[…]
8:return $\hat{\mathbf{\Phi}}$, $\mathbb{D}_{\text{GW}}$